Swifty Corporation has 4,100 shares of 9%, $102 par value preferred stock outstanding at December 31, 2017. At December 31, 2017, the company declared a $127,500 cash dividend.

Determine the dividend paid to preferred stockholders and common stockholders under each of the following scenarios.

1. The preferred stock is noncumulative, and the company has not missed any dividends in previous years.

The dividend paid to preferred stockholders

$enter a dollar amount

The dividend paid to common stockholders

$enter a dollar amount


2. The preferred stock is noncumulative, and the company did not pay a dividend in each of the two previous years.

The dividend paid to preferred stockholders

$enter a dollar amount

The dividend paid to common stockholders

$enter a dollar amount


3. The preferred stock is cumulative, and the company did not pay a dividend in each of the two previous years.

The dividend paid to preferred stockholders

$enter a dollar amount

The dividend paid to common stockholders

$enter a dollar amount
Answer #1

Preferred dividend= 4100*102*9% = 37638

1)

The dividend paid to preferred stockholders

$ 37638

The dividend paid to common stockholders

(127500-37638)= $ 89862

2)

The dividend paid to preferred stockholders

$ 37638

The dividend paid to common stockholders

(127500-37638)= $ 89862

3)

The dividend paid to preferred stockholders

(37638*3)= 112914

The dividend paid to common stockholders

(127500-112914)= $ 14586